OpenStudy (anonymous):

how to get equation o. line in slope-intercept form through (-1,-1) and parallel to y= -x -5

OpenStudy (anonymous):

when 2 lines are parallel the slopes are equal, and the slope of y=-x-5 is m= -1 from y=mx+c eqn. and line through point (-1,-1) is (y-y1)=m(x-x1) => y=-x is resultant eqn of line

OpenStudy (akashdeepdeb):

Is this doubt cleared @angelsmg22 ?

OpenStudy (anonymous):

No, a bit more of an explanation would help

OpenStudy (akashdeepdeb):

Okay look. Lets go step by step. :D

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (akashdeepdeb):

STEP 1: Parallel lines are almost the same lines but just beside each other. So we say that they have the same slope. Understood this? :D

OpenStudy (anonymous):

Ok, yes

OpenStudy (akashdeepdeb):

STEP 2: So this new parallel line has the same slope and passes through the point (-1,-1) Okay?

OpenStudy (anonymous):

Yes

OpenStudy (akashdeepdeb):

STEP 3: Now do you know how to find the slope of a line? 1) RISE/RUN 2) From a line's equation?

OpenStudy (anonymous):

Ok, yes

OpenStudy (akashdeepdeb):

Then find the slope of the line whose equation is already given and tell me! :D

OpenStudy (anonymous):

-1

OpenStudy (akashdeepdeb):

Excellent. So we can infer from this that the slope of the parallel line would be -1 and would pass through (-1,-1). Right? :D

OpenStudy (anonymous):

yes

OpenStudy (akashdeepdeb):

STEP 4/5 [I lost track] Now do you know the slope-intercept form of a line? :D

OpenStudy (anonymous):

no..

OpenStudy (akashdeepdeb):

Okay it is. y-y1 = m(x-x1) When m= -1 (x1,y1) = (-1,-1) Understood? :D

OpenStudy (anonymous):

yes, thank you !

OpenStudy (akashdeepdeb):

This is what @ramyav07 basically said in a very short and concise manner. :)